Camp layout and skiway taking shape – and we had visitors – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

NEEM > Field diaries folder > UK diaries 2015 > Camp layout and skiway...

28 May 2015

Camp layout and skiway taking shape – and we had visitors

A group of happy people after the traverse to EGRIP. From left lower row: Mathias, Sverrir, Dorthe, Paul, Helle, Nanna and Anna. From left upper row: J.P., Sepp, Joel and Pat.

A group of happy people after the traverse to EGRIP. From left lower row: Mathias, Sverrir, Dorthe, Paul, Helle, Nanna and Anna. From left upper row: J.P., Sepp, Joel and Pat.



Survey of camp area is now complete, we can now recognize the future layout of camp.

As skiway marking is approaching the end, we now have a significant structure to look at for reference.

We got a visit from 11.00 to 19.00 by a group of scientists that set-up an automatic station at EGRIP. They arrived from Summit and they returned to Summit. The twin Otter inaugurated the skiway as the snow at EGRIP is bumpy, so they took advantage of a freshly groomed skiway to take off. The visitors had lunch at EGRIP; but otherwise they focussed on getting their job done.

It was a day where the science groups had a little time to do sampling and experiments. Tomorrow, we will begin to rebuild the garages that were taken down at NEEM.

Work is progressing fast, and we have almost caught up with the six days delay in departure from NEEM.

If weather continues to behave, we will be ready to receive the first LC-130 flight on Monday.

Our Field Operations Management (FOM) office in Kangerlussuaq is open again, as Christine Hvidberg arrived from Copenhagen today. The EGRIP Field Leader forgot all about normal telephone hours, as it has been a while since the FOM office last was open.

What we have done today:

1.    Finished marking approaches and apron.
2.    Unpacking sleds and arranging cargo in cargo line.
3.    Building berm for future carpenters garage and foundation for dome.
4.    Visit by Twin Otter from the Abdalati group. The group set up an automatic
       station.
5.    Drilled shallow core.
6.    Performed pit studies of surface snow.
7.    Activated British Antarctic Survey, Hugh Corr, radar.
8.    Groomed skiway.

Weather: Beautiful day, some clouds and a tendency for fog. Temp. – 12°C to -27°C, 0-3 m/s from SW and W. Visibility:  Un-restricted, down to 1 km in fog banks.

FL, J.P. Steffensen


← Previous entry   Next entry →